### v2

@@ 3.1 Local Linear Regressions
$\widehat{\tau}_{\texttt{llr}}(x^{c};h):=$ $\widehat{\mu}_{\texttt{llr}}^{+}(x^{c};h)-\widehat{\mu}_{\texttt{llr}}^{-}(x^{c};h),$ (2)
@@
with respect to $h$. However, Calonico et al. (2014) note that the MSE-optimal bandwidth obtained from optimizing (5), denoted as $h^{*}$, is oversmoothing in the sense that $n(h^{{*}})^{5}$ does not converge to 0, leading to a bias term that must be corrected in order to conduct valid inference. They propose subtracting from $\widehat{\tau}_{\texttt{llr}}(x^{c};h^{*})$ an estimated bias term and then adjust the standard error to account for the variation in the estimation of the bias term. This procedure, however, does not readily generalize to multivariate RD designs. Although Zajonc (2012) proposes a similar MSE-optimal bandwidth selector in the bivariate case, valid inference still requires bias correction due to oversmoothing. In empirical research involving multiple scores, the multivariate problem is usually reduced to one dimension so that univariate methods can apply.
3.1.1 Problem with Transforming Multivariate Scores
As mentioned earlier, common practices to reduce a multivariate RD design to a univariate one include subsetting the data and transforming the multivariate score to a one-dimensional running variable, and each has its own pitfalls. In particular, transforming a multivariate score could incur unexpected behaviors of the resulting density of the effective univariate running variable. In this section, we detail a situation that may arise under such transformation: the transformed univariate score may have zero density at the new cutoff, violating a key assumption for both the identification and estimation of treatment effects in RD designs.
In Section 4, we use real data from Keele and Titiunik (2015) to generate simulation data for the Monte Carlo study. Keele and Titiunik (2015) transform a bivariate score to a univariate one using rescaled Euclidean (chordal) distance to a given treatment boundary point and then estimate the treatment effect at the new cutoff $x^{c}=0$ using local linear regression with robust bias correction as in Calonico et al. (2014). The following proposition gives a general characterization of bivariate densities that have a zero density at the effective univariate cutoff $x^{c}=0$, once the bivariate score is collapsed to a univariate variable by taking the Euclidean distance between the bivariate score and a given boundary point.
@@ 3.1.1 Problem with Transforming Multivariate Scores
where $l(e)$ and $u(e)$ are the lower and upper bound of $v$ that can depend on $e$ [^3] To simplify notation, we assume without loss of generality that the lower and upper bounds of $(e,v)$ are attained so that the joint support of $(E,V)$ is a closed subset of $\mathbb{R}^{2}$; otherwise the closed brackets need to be changed to open brackets accordingly. Then the Euclidean distance $E$ has a zero density at $e=0$, $f_{E}(0)=0$, if and only if
@@
We give the proof of Proposition 1 and two examples under which (6) holds in Appendix A. This characterization depends on the population joint density of the scores and is specific to the transformation chosen by the researcher. To illustrate, the first panel of Figure 1 is the true density of a transformed bivariate score that is uniformly distributed over $[-1,1]^{2}$, and the point to which we calculate the Euclidean distance is $(0,0)$; see Example A.1 for the derivation of this density. As in Keele and Titiunik (2015), we sign the Euclidean distance so that treated (control) observations receive a positive (negative) transformed score, and the new cutoff after transformation becomes 00, at which the density of the transformed score is exactly 00. In practice, the researcher has access to only finite samples, and problems of this kind are difficult to detect without knowing the true DGP. In the last two panels of Figure 1, we plot the raw histograms of the transformed bivariate score with respect to a given treatment boundary point using the exact same transformation employed by Keele and Titiunik (2015); each histogram uses data simulated from a different DGP built using the real data from Keele and Titiunik (2015), and thus the simulated bivariate scores do not follow any known distributions. [^4] We provide details on how the simulation data is generated in Section 4. In all three panels, the (empirical) density of the transformed bivariate score is zero at the effective univariate cutoff $x^{c}=0$. When such zero-density problem arises, not only is the conditional treatment effect at the transformed cutoff not identified because the conditional expectations are not defined, the problem is also reflected in estimation, especially for very large sample sizes, as we show in the simulations in Section 4.
Figure 1: True and empirical densities of transformed bivariate scores. From left to right: (i) True density of a bivariate uniform running variable on $[-1,1]^{2}$ transformed using signed Euclidean distance; (ii) histogram of transformed bivariate coordinates using signed chordal distance to the selected boundary point, plotted using all $40$ million observations drawn from the WGAN-generated DGP; (iii) histogram of transformed bivariate coordinates using signed chordal distance to the selected boundary point, plotted using a sample of $1$ million observations from the polynomial DGP. The number of bins for both histograms is 10,000.
@@ 3.2 A Minimax-Optimal Estimator
To avoid the problem of choosing a multivariate bandwidth, Imbens and Wager (2019) propose an estimator based on numerical optimization that is minimax-optimal among all estimators that are linear in the outcome variable over all problems with $\sigma_{i}^{2}:=Var[Y_{i}|X_{i}]$ and a second derivative bound $\mathcal{B}$ on the CEF of potential outcomes. Specifically, their estimator for $\tau(x^{c})$ takes the form
@@
$\widehat{\gamma}(x^{c};\mathcal{B})$ $=\arg\min_{\gamma}\left\{\sum_{i=1}^{n}\gamma_{i}^{2}\sigma_{i}^{2}+b^{2}(\gamma;\mathcal{B})\right\},$ (8)
@@
$\mathcal{I}_{\alpha}:\liminf_{n\to\infty}\inf\left\{\mathbb{P}[\tau(x^{c})\in\mathcal{I}_{\alpha}]:\forall\,x,\lVert\nabla^{2}\mu^{+}(x)\rVert\leq\mathcal{B}\text{ and }\lVert\nabla^{2}\mu^{-}(x)\rVert\leq\mathcal{B}\right\}\geq 1-\alpha.$ (9)
Implementing the minimax-optimal estimator (7) requires estimates of $\sigma_{i}^{2}$ and $\mathcal{B}$. Imbens and Wager (2019) recommend estimating $\sigma_{i}^{2}$ by averaging the residuals from an ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $A(X_{i})$. Estimating $\mathcal{B}$, however, is problem-specific and less straightforward. Imbens and Wager (2019, p.273) recommend estimating the CEF globally using, for example, second-order polynomials, and then multiplying the maximal estimated curvature by 2 or 4. In one of our simulations in Section 4 using data from Lee (2008), we find that using second-order polynomials yields estimates of $\mathcal{B}$ that are too small, and multiplying by $4$ does not help; only after multiplying the maximal estimated curvature by $30$ do we get an estimate that is close to the true (least upper) bound $\mathcal{B}$.
@@ 3.3 Honest Regression Forests and Local Linear Forests
We now introduce two additional theoretically valid estimators that can be leveraged in multivariate RD designs but have been largely overlooked in the literature—the honest regression forests of Wager and Athey (2018) and the local linear forests of Friedberg et al. (2020). Both estimators are variants of random forests originally proposed by Breiman (2001) and have received popularity in estimating heterogeneous treatment effects under unconfoundedness.
@@ 3.3.1 Honest Regression Forests
A regression forest evaluated at some point $x\in\mathcal{X}$, denoted as $RF(x)$, is an average of $B$ trees, $\{T_{b}(x)\}_{b=1}^{B}$, each fitted using a bootstrapped sample or a random subsample of $\mathcal{S}_{n}:=\{(Y_{i},X_{i})\}_{i=1}^{n}$. Following Athey et al. (2019), Friedberg et al. (2020), and Wager and Athey (2018), we focus on forests built using random subsamples, for which central limit theorems are established. Given a subsample of size $s$ from $\mathcal{S}_{n}$, $\{(Y_{i_{k}},X_{i_{k}})\}_{k=1}^{s}$, a regression tree recursively divides the sample into mutually exclusive and collectively exhaustive convex subsets called leaves. For a point $x\in\mathcal{X}$, let $L(x)$ denote the leaf that contains $x$. A tree $T$ estimates the CEF, $\mathbb{E}[Y_{i}|X_{i}=x]$, by averaging the observations that fall into the same leaf as $x$:
@@
$\mathbb{E}\left[\bigl{(}Y_{i}-T(X_{i})\bigr{)}^{2}\right],$ (10)
and searching for all possible partitions is computationally infeasible. Instead, standard splitting rules rely on greedy algorithms that immediately improve the in-sample fit of leaf-wide averages after each split.
Estimates from a single tree can be highly unstable without regularization. In the extreme case, one can achieve perfect in-sample fit by partitioning the sample data so that each leaf contains only one observation. Regression forests stabilize predictions from individual trees by averaging many of them: for a collection of $b=1,...,B$ trees, we fit each tree $T_{b}(x)$ using a random subsample of size $s$, and obtain the regression forest estimate by averaging across all $B$ trees, $\frac{1}{B}\sum_{b=1}^{B}T_{b}(x)$. To achieve a sizable reduction in variance, we may want to decorrelate the trees as much as possible by, for example, randomly selecting a dimension to split at each node. We assume throughout that $B$ is large enough so that the Monte Carlo variability of which subsamples are drawn is negligible and the following definition of regression forests from Wager and Athey (2018) given the training sample $\mathcal{S}_{n}$ and subsample size $s$ applies:
$RF\bigl{(}x;s,\mathcal{S}_{n}\bigr{)}:={\binom{n}{s}}^{-1}\sum_{1\leq i_{1}<...<i_{s}\leq n}\mathbb{E}_{\xi}\bigl{[}T(x;\xi,\{(Y_{i_{k}},X_{i_{k}})\}_{k=1}^{s})\bigr{]},$ (11)
@@
$\widehat{\tau}_{\texttt{rf}}(x^{c})$ $:=\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})-\widehat{\mu}_{\texttt{rf}}^{-}(x^{c}),$ (12)
@@
$\widehat{\mu}_{\texttt{rf}}^{-}(x^{c})$ $:=RF\bigl{(}x^{c};s_{0},\{(Y_{i},X_{i})\in\mathcal{S}_{n}:A(X_{i})=0\}\bigr{)}.$
@@
Remark 1. Although it seems that the local linear estimator (2) depends on one more parameter, $h$, than the regression forest estimator (12), we note that this is not the case. Each tree $T$ in the regression forest implicitly defines a neighborhood $L(x^{c})$ for the treatment boundary point $x^{c}$. These neighborhoods are then aggregated, producing a weight for each observation that is proportional to the fraction of time it falls into the same leaf as $x^{c}$ across all trees in the regression forest. These forest weights are analogous to the weights generated by a kernel with bandwidth $h$. Thus (12) implicitly depends on some “pseudo-bandwidth” generated by tree partitions across the forest. Analogously, the $\gamma$ weights of the minimax-optimal estimator (7) play the role of kernels and bandwidths.
We can equivalently express regression forests as solutions to weighted least-squares constant regressions:
@@
where $L^{+}_{b}(x^{c})$ denotes the leaf containing $x^{c}$ in the $b$-th tree fitted using the treated observations; $W_{\texttt{rf}}^{-}$ is defined similarly for the control observations. More generally, the least-square representation of forests in (13) and (14) effectively views $\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})$ and $\widehat{\mu}_{\texttt{rf}}^{-}(x^{c})$ as solutions to the local moment conditions,
$\mathbb{E}\left[(Y_{i}-\theta)^{2}|X_{i}=x^{c},D_{i}=1\right]=0\quad\text{and}\quad\mathbb{E}\left[(Y_{i}-\theta)^{2}|X_{i}=x^{c},D_{i}=0\right]=0,$
and solves for $\theta$ using their empirical analogues with forest weights that reflect the importance of each observation $i$ in terms of minimizing the objective function. Athey et al. (2019) use forests as weighting kernels to solve parameters identified by local moments in a more general setting.
@@ 3.3.2 Honest Local Linear Forests
$\widehat{\tau}_{\texttt{llf}}(x^{c})$ $:=\widehat{\mu}_{\texttt{llf}}^{+}(x^{c})-\widehat{\mu}_{\texttt{llf}}^{-}(x^{c}),$
$\widehat{\mu}_{\texttt{llf}}^{+}(x^{c})$ $:=\bm{e}_{1}^{\intercal}\left(\arg\min_{(\beta_{0},\beta_{1})^{\intercal}}\sum_{i=1}^{n}D_{i}W_{\texttt{rf}}^{+}(X_{i},x^{c})\bigl{(}Y_{i}-\beta_{0}-\beta_{1}^{\intercal}(X_{i}-x^{c})\bigr{)}^{2}+\lambda\|\beta_{1}\|_{2}^{2}\right),$ (16)
$\widehat{\mu}_{\texttt{llf}}^{-}(x^{c})$ $:=\bm{e}_{1}^{\intercal}\left(\arg\min_{(\beta_{0},\beta_{1})^{\intercal}}\sum_{i=1}^{n}(1-D_{i})W_{\texttt{rf}}^{-}(X_{i},x^{c})\bigl{(}Y_{i}-\beta_{0}-\beta_{1}^{\intercal}(X_{i}-x^{c})\bigr{)}^{2}+\lambda\|\beta_{1}\|_{2}^{2}\right),$ (17)
@@
The assumptions needed for the asymptotic normality of local linear forests overlap to a large extent with the conditions required for honest regression forests to be asymptotically normal. However, with an additional assumption on the smoothness of the CEF $\mathbb{E}[Y_{i}|X_{i}=x]$—that it is differentiable with a Lipschitz continuous derivative, as opposed to the Wager and Athey (2018) condition that only requires it to be Lipschitz continuous—Friedberg et al. (2020) show an improved convergence rate of local linear forests relative to the rate of regression forests, effectively exploiting the additional smoothness assumption with a linear adjustment. We refer the reader to Wager and Athey (2018) and Friedberg et al. (2020) for details on the technical assumptions and discussions thereof. Pointwise confidence intervals for both honest regression forests and honest local linear forests are constructed using variance estimates from the bootstrap of little bags algorithm of Sexton and Laake (2009), with consistency results established by the random forest delta method of Athey et al. (2019, Theorem 6). In our Monte Carlo study, we find that local linear forests are in general less biased than regression forests and have better coverage rates whenever the simulation DGP exhibits strong linear trends.
@@
Remark 2. One caveat of using local linear forests in multivariate RD designs is that the point at which we evaluate local linear forests, $x^{c}\in\mathscr{B}$, is necessarily a boundary point of the support of $X_{i}$ for both treated and control populations. On the other hand, the central limit theorem of Friedberg et al. (2020, Theorem 1) only applies to points in the interior of the support, as the proof starts with a usual Taylor expansion of $\mathbb{E}[Y_{i}|X_{i}=x]$ around $x^{c}$. [^6] Asymptotic normality of local linear regressions also relies on a similar Taylor expansion, but it is the CEF of potential outcomes that gets expanded (as opposed to the CEF of observed outcomes conditional on either the treated or the control population), and so in this case $x^{c}\in\mathscr{B}$ is an interior point of the support of $X$. We sidestep this problem by adding an arbitrarily small “buffer” to $x^{c}\in\mathscr{B}$: we pick another point $\tilde{x}^{c}$ in the interior of the treated score region such that $\|\tilde{x}^{c}-x^{c}\|=\epsilon$ for an arbitrarily small $\epsilon>0$ and evaluate the treated local linear forest $\widehat{\mu}^{+}_{\texttt{llf}}$ at $\tilde{x}^{c}$ to approximate $\widehat{\mu}^{+}_{\texttt{llf}}(x^{c})$; we do the same for the control local linear forest $\widehat{\mu}^{-}_{\texttt{llf}}$. We require $\mu^{+}(x^{c})$ and $\mu^{-}(x^{c})$ to satisfy the assumptions in Friedberg et al. (2020, Theorem 1), which imply that both $\mu^{+}(x^{c})$ and $\mu^{-}(x^{c})$ are Lipschitz in $x^{c}$ and therefore very close to $\mu^{+}(\tilde{x}^{c})$ and $\mu^{-}(\tilde{x}^{c})$, respectively, for small $\epsilon$. Ideally, one should adjust the confidence interval accordingly to account for such approximation, but in our simulations we choose $\epsilon=10^{-30}$ so that such correction is negligible up to machine precision.

### v3

@@ 3.1 Local Linear Regressions
$\widehat{\tau}_{\texttt{llr}}(x^{c};h):=\widehat{\mu}_{\texttt{llr}}^{+}(x^{c};h)-\widehat{\mu}_{\texttt{llr}}^{-}(x^{c};h).$ (3)
@@
with respect to $h$. However, Calonico et al. (2014) note that the MSE-optimal bandwidth obtained from optimizing (4), denoted as $h^{*}$, is oversmoothing in the sense that $n(h^{{*}})^{5}$ does not converge to 0, leading to a bias term that must be corrected in order to conduct valid inference. They propose subtracting from $\widehat{\tau}_{\texttt{llr}}(x^{c};h^{*})$ an estimated bias term and then adjusting the standard error to account for the uncertainty in the bias estimation. This procedure, however, does not readily generalize to multivariate RD designs. Although Zajonc (2012) proposes a similar MSE-optimal bandwidth selector in the bivariate case, valid inference still requires bias correction due to oversmoothing, which is technically challenging beyond one dimension and to our knowledge there is no such proposal in the existing literature. In empirical research involving multiple scores, the multivariate problem is usually reduced to one dimension so that univariate methods can apply.
3.1.1 A Problem with Transforming Multivariate Scores
As mentioned earlier, common practices of reducing a multivariate RD design to a univariate one include subsetting the data and reducing the multivariate score to one dimension, and each has its own pitfalls. In particular, transforming a multivariate score in some cases incurs unexpected behaviors of the resulting density of the effective univariate running variable. In this section, we detail a situation that may arise under such transformation: the transformed univariate score may have zero density at the new cutoff, violating a key assumption for both the identification and estimation of treatment effects in RD designs.
We illustrate this problem for a transformation used by Keele and Titiunik (2015), who transform a bivariate score to a univariate one using rescaled Euclidean (chordal) distance to a given treatment boundary point and then estimate the treatment effect at the new cutoff $x^{c}=0$ using local linear regression with robust bias correction as in Calonico et al. (2014). The following proposition gives a general characterization of bivariate densities that have a zero density at the effective univariate cutoff $x^{c}=0$, once the bivariate score is collapsed to a univariate variable by taking the Euclidean distance between the bivariate score and a given boundary point.
@@ 3.1.1 A Problem with Transforming Multivariate Scores
where $l(e)$ and $u(e)$ are the lower and upper bound of $v$ that can depend on $e$ [^4] To simplify notation, we assume without loss of generality that the lower and upper bounds of $(e,v)$ are attained so that the joint support of $(E,V)$ is a closed subset of $\mathbb{R}^{2}$; otherwise the closed brackets need to be changed to open brackets accordingly. Then the Euclidean distance $E$ to the given point $(c_{1},c_{2})$ has zero density at the new cutoff $e=0$, $f_{E}(0)=0$, if and only if
@@
Intuitively, such transformation involves a loss of information: many distinct multivariate realizations can map to the same distance on the real line. In particular, for the collapsed variable to be zero, all coordinates of the original variable must simultaneously have zero distance, which is a rare event. We give the proof of Proposition 1 and two examples under which (5) holds in Appendix A. This characterization depends on the population joint density of the scores and is specific to the transformation chosen by the researcher.
Figure 1: True and empirical densities of transformed bivariate scores. From left to right: (i) True density of a bivariate uniform running variable on $[-1,1]^{2}$ transformed by signed Euclidean distance; (ii) histogram of transformed bivariate coordinates using signed chordal distance to the selected boundary point, plotted using all $40$ million observations drawn from the WGAN-generated DGP; (iii) histogram of transformed bivariate coordinates using signed chordal distance to the selected boundary point, plotted using a sample of $1$ million observations from the polynomial DGP. The number of bins for both histograms is 10,000.
To illustrate, the first panel of Figure 1 is the true density of a transformed bivariate score that is uniformly distributed over $[-1,1]^{2}$, and the point to which we calculate the Euclidean distance is $(0,0)$; see Example A.1 for the derivation of this density. As in Keele and Titiunik (2015), we sign the Euclidean distance so that treated (control) observations receive a positive (negative) transformed score, and the new cutoff after transformation becomes 00, at which the density of the transformed score is exactly 00. In practice, the researcher has access to only finite samples, and problems of this kind are difficult to detect without knowing the true DGP. In the last two panels of Figure 1, we plot the raw histograms of the transformed bivariate score with respect to a given treatment boundary point using the exact same transformation employed by Keele and Titiunik (2015); each histogram uses data simulated from a different DGP built using the real data from Keele and Titiunik (2015), and thus the simulated bivariate scores do not follow any known distributions. [^5] We provide details on how the simulation data is generated in Section 4. In all three panels, the empirical density of the transformed bivariate score, be it true or empirical, is zero at the effective univariate cutoff $x^{c}=0$. When such zero-density problem arises, not only is the conditional treatment effect at the transformed cutoff not identified because the conditional expectations are not defined, the problem is also reflected in estimation, especially for very large sample sizes, as we show in the simulations in Section 4.
@@ 3.2 A Minimax-Optimal Estimator
To avoid the problem of choosing a multivariate bandwidth, Imbens and Wager (2019) propose an estimator based on numerical optimization that is minimax-optimal among all estimators that are linear in the outcome variable over all problems with finite $\sigma_{i}^{2}:=Var[Y_{i}|X_{i}]$ and a second derivative bound $\mathcal{B}<\infty$ on the CEF of potential outcomes. Specifically, their estimator for $\tau(x^{c})$ takes the form
@@
$\widehat{\gamma}(x^{c};\mathcal{B})=\arg\min_{\gamma\in\mathbb{R}^{n}}\left\{\sum_{i=1}^{n}\gamma_{i}^{2}\sigma_{i}^{2}+b^{2}(\gamma;\mathcal{B})\right\},$
@@
$\mathcal{I}_{\alpha}:\liminf_{n\to\infty}\inf\left\{\mathbb{P}[\tau(x^{c})\in\mathcal{I}_{\alpha}]:\forall\,x,\lVert\nabla^{2}\mu^{+}(x)\rVert\leq\mathcal{B}\text{ and }\lVert\nabla^{2}\mu^{-}(x)\rVert\leq\mathcal{B}\right\}\geq 1-\alpha.$
Implementing the minimax-optimal estimator (6) requires estimates of $\sigma_{i}^{2}$ and $\mathcal{B}$. Imbens and Wager (2019) recommend estimating $\sigma_{i}^{2}$ by averaging the residuals from an ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $A(X_{i})$. Estimating $\mathcal{B}$, however, is problem-specific and less straightforward. Imbens and Wager (2019, p.273) recommend estimating the CEFs globally using second-order polynomials and then multiplying the maximal estimated curvature by 2 or 4. In one of our simulations in Section 4 using data from Lee (2008), we find that using second-order polynomials yields estimates of $\mathcal{B}$ that are too small, even after multiplying by $4$; only after multiplying the maximal estimated curvature by $30$ do we get an estimate that is close to the true (least upper) bound $\mathcal{B}$.
@@ 3.3 Honest Regression Forests and Local Linear Forests
We now introduce two additional theoretically valid estimators that can be leveraged in multivariate RD designs but have been largely overlooked in the literature—the honest regression forests of Wager and Athey (2018) and the local linear forests of Friedberg et al. (2020). Both are variants of random forests originally proposed by Breiman (2001) and have received popularity in estimating heterogeneous treatment effects under unconfoundedness.
@@ 3.3.1 Honest Regression Forests
A regression forest evaluated at some point $x\in\mathcal{X}$, denoted as $RF(x)$, is an average of $B$ trees, $\{T_{b}(x)\}_{b=1}^{B}$, each fitted using a bootstrapped sample or a random subsample of $\mathcal{S}_{n}:=\{(Y_{i},X_{i})\}_{i=1}^{n}$. Following Athey et al. (2019), Friedberg et al. (2020), and Wager and Athey (2018), we focus on forests built using random subsamples, for which central limit theorems are established. Given a subsample of size $s$ from $\mathcal{S}_{n}$, $\{(Y_{i_{k}},X_{i_{k}})\}_{k=1}^{s}$, a regression tree recursively divides the sample into mutually exclusive and collectively exhaustive convex subsets called leaves. For a point $x\in\mathcal{X}$, let $L(x)$ denote the leaf that contains $x$. A tree $T$ estimates the CEF, $\mathbb{E}[Y_{i}|X_{i}=x]$, by averaging the outcomes of observations that fall into the same leaf as $x$:
@@
$\mathbb{E}\left[\bigl{(}Y_{i}-T(X_{i})\bigr{)}^{2}\right],$
and searching for all possible partitions is computationally infeasible. Instead, standard splitting rules rely on greedy algorithms that immediately improve the in-sample fit of leaf-wide averages after each split. As such, estimates from a single tree can be highly unstable without regularization: in the extreme case, one can achieve perfect in-sample fit by partitioning the sample data so that each leaf contains only one observation.
Regression forests stabilize predictions from individual trees by averaging many of them: for a collection of $b=1,...,B$ trees, we fit each tree $T_{b}(\cdot)$ using a random subsample of size $s$, and obtain the regression forest estimate by averaging the predictions across all $B$ trees, $\frac{1}{B}\sum_{b=1}^{B}T_{b}(x)$. To achieve a sizable reduction in variance, we may want to decorrelate the trees as much as possible by, for example, randomly selecting a dimension to split at each node. Following Wager and Athey (2018), we assume throughout that $B$ is large enough so that the Monte Carlo variability of which subsamples are drawn is negligible and adopt their definition of regression forests given the training sample $\mathcal{S}_{n}$ and subsample size $s$:
$RF\bigl{(}x;s,\mathcal{S}_{n}\bigr{)}:={\binom{n}{s}}^{-1}\sum_{1\leq i_{1}<...<i_{s}\leq n}\mathbb{E}_{\xi}\bigl{[}T(x;\xi,\{(Y_{i_{k}},X_{i_{k}})\}_{k=1}^{s})\bigr{]},$
@@
$\widehat{\mu}_{\texttt{rf}}^{-}(x^{c})$ $:=RF\bigl{(}x^{c};s_{0},\{(Y_{i},X_{i})\in\mathcal{S}_{n}:A(X_{i})=0\}\bigr{)},$
$\widehat{\tau}_{\texttt{rf}}(x^{c})$ $:=\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})-\widehat{\mu}_{\texttt{rf}}^{-}(x^{c}).$ (7)
@@
Remark 1. Although it seems that the local linear estimator (3) depends on one more parameter, $h$, than the regression forest estimator (7), we note that this is not the case. Each tree $T$ in the regression forest implicitly defines a neighborhood $L(x^{c})$ for the treatment boundary point $x^{c}$. These neighborhoods are then aggregated, producing a weight for each observation that is proportional to the fraction of time it falls into the same leaf as $x^{c}$ across all trees in the regression forest. These forest weights are analogous to the weights generated by a kernel with bandwidth $h$. Thus (7) implicitly depends on some “pseudo-bandwidth” generated by tree partitions across the forest. Analogously, the $\gamma$ weights of the minimax-optimal estimator (6) play the role of kernels and bandwidths. We can equivalently express regression forests as solutions to weighted least-squares constant regressions:
@@
where $L^{+}_{b}(x^{c})$ denotes the leaf containing $x^{c}$ in the $b$-th tree fitted using the treated observations; $W_{\texttt{rf}}^{-}$ is defined similarly for the control observations. More generally, the least-square representation of forests in (8) and (9) effectively views $\widehat{\mu}_{\texttt{rf}}^{+}(x^{c})$ and $\widehat{\mu}_{\texttt{rf}}^{-}(x^{c})$ as solutions to $\mathbb{E}\left[(Y_{i}-\theta)^{2}|X_{i}=x^{c},D_{i}=1\right]=0$ and $\mathbb{E}\left[(Y_{i}-\theta)^{2}|X_{i}=x^{c},D_{i}=0\right]=0$, and solves for $\theta$ using their empirical analogs with forest weights reflective of the importance of each observation $i$ in terms of minimizing the objective function. Athey et al. (2019) use forests as weighting kernels to solve parameters identified by local moments in a more general setting.
@@ 3.3.2 Honest Local Linear Forests
[…]
@@
The assumptions needed for the asymptotic normality of local linear forests largely overlap with those required for honest regression forests to be asymptotically normal. However, with an additional assumption on the smoothness of the CEF $\mathbb{E}[Y_{i}|X_{i}=x]$—that it is differentiable with a Lipschitz continuous derivative, as opposed to the Wager and Athey (2018) condition that only requires it to be Lipschitz continuous—Friedberg et al. (2020) show an improved convergence rate of local linear forests relative to the rate of regression forests, effectively exploiting the additional smoothness assumption with a linear adjustment. We refer the reader to Wager and Athey (2018) and Friedberg et al. (2020) for details on the technical assumptions and discussions thereof. Pointwise confidence intervals for both honest regression forests and honest local linear forests are constructed using variance estimates from the bootstrap of little bags algorithm of Sexton and Laake (2009), with consistency results established by the random forest delta method of Athey et al. (2019, Theorem 6). In our Monte Carlo study, we find that local linear forests are in general less biased than regression forests and have better coverage rates whenever the simulation DGP exhibits strong linear trends.
@@
Remark 2. One caveat of using local linear forests in multivariate RD designs is that the point at which we evaluate local linear forests, $x^{c}\in\mathscr{B}$, is necessarily a boundary point of the support of $X_{i}$ for both treated and control populations. On the other hand, the central limit theorem of Friedberg et al. (2020, Theorem 1) only applies to points in the interior of the support, as the proof starts with a usual Taylor expansion of $\mathbb{E}[Y_{i}|X_{i}=x]$ around $x^{c}$. [^7] Asymptotic normality of local linear regressions also relies on a similar Taylor expansion, but it is the CEFs of potential outcomes that get expanded (as opposed to the CEFs of observed outcomes of either the treated or the control population), and so in this case $x^{c}\in\mathscr{B}$ is an interior point of the support of $X$. We sidestep this problem by adding an arbitrarily small “buffer” to $x^{c}\in\mathscr{B}$: we pick another point $\tilde{x}^{c}$ in the interior of the treated score region such that $\|\tilde{x}^{c}-x^{c}\|=\epsilon$ for an infinitesimal $\epsilon>0$ and evaluate the treated local linear forest $\widehat{\mu}^{+}_{\texttt{llf}}$ at $\tilde{x}^{c}$ to approximate $\widehat{\mu}^{+}_{\texttt{llf}}(x^{c})$; we do the same for the control local linear forest $\widehat{\mu}^{-}_{\texttt{llf}}$. Under the assumptions in Friedberg et al. (2020, Theorem 1), both $\mu^{+}(x^{c})$ and $\mu^{-}(x^{c})$ are Lipschitz in $x^{c}$ and therefore very close to $\mu^{+}(\tilde{x}^{c})$ and $\mu^{-}(\tilde{x}^{c})$, respectively, for an infinitesimal $\epsilon$. Ideally, one should adjust the confidence interval accordingly to account for such approximation. We leave this for future research, and in our simulations and empirical application we choose $\epsilon=10^{-30}$ so that such correction is essentially negligible up to machine precision.